Three resistors A, B, and C are connected in series to a 120 volt source. If RA is = 64 ohms and EB = 40volts, when the current is 0.5 amp, calculate the resistances RB and RC.

1
Expert's answer
2022-01-05T14:40:04-0500

The equivalent resistance of the three resistors is the source voltage over the current:


"R_\text{eq}=R_A+R_B+R_C=E/I=120/0.5=240\Omega.\\
R_B=E_B/I=40/0.5=80\Omega.\\
R_C=R_\text{eq}-R_A-R_B=240-64-80=96\Omega."